Suppose that the quantity of steel demanded in France is given by QS = 100 − 2PS + 0.5Y + 0.2PA, where QS is the quantity of steel demanded per year, PS is the market price of steel, Y is real GDP in France, and PA is the market price of aluminum. In 2011, PS = 10, Y = 40, and PA = 100. How much steel will be demanded in 2011? What is the price elasticity of demand, given market conditions in 2011?
Answer #1

Answer : By putting given values in demand function we get,

Qs = 100 - (2 * 10) + (0.5 * 40) + (0.2 * 100)

=> Qs = 100 - 20 + 20 + 20

=> Qs = 120

Therefore, in 2011 the quantity demanded for steel will be Qs = 120.

Price elasticity of demand (Es) = (Qs / Ps) * (Ps / Qs)

=> Es = - 2 * (10 / 120) = - 2 * 0.08

=> Es = - 0.2

So, in 2011 the price elasticity of demand is, Es = - 0.2 .